OpenStudy (anonymous):

solve the system of equations by graphing then classify the system as consistent or inconsistent and as dependent or independent 5x-5y=-15 5y-5x=15

OpenStudy (anonymous):

graph and you will see they are the same line

OpenStudy (anonymous):

Without graphing, you can look at it and tell which it is.\[5x-5y=-15\]\[-1(5y-5x)=(-1)(15)\]\[5y-5x=15\]which is the same equation as your second line...so they are the same graph.